Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
amagazzu - 0:36 Friday 24 March 2023 (59399) Print this report
Operator Report - Afternoon shift

The shift was dedicated to the planned ISC activity, CARM2RFC & LSC noise, carried out by Pinto and Boldrini.
The ITF unlocked from LOW_NOISE_1 at 16:29 UTC, taking advantage of the event Gosselin went to the WE Building to investigate the PSTAB (see report #59397), while Allocca went to the NE Building to work on the Tiltometer. Both activites were concluded at around 17:00 UTC.
After that we reached CARM_NULL_1F and LOW_NOISE_1 multiple times, but the lock lasted only around 30 minutes. At around 19:15 UTC Boldrini noted the loop of B1p QD1 galvo was not working anymore, so we contacted the DET Oncall to solve the issue. The problem was solved at around 20:40 UTC and we continued the rest of the shift with the support of Gouaty.
ITF left in CARM_NULL_1F, TCS INJ measurement in progress.

Oncall events

DET
(23-03-2023 19:15 - 23-03-2023 20:40) From remote
Status: Ended
Description: At around 19:15 UTC Boldrini noted the that B1p QD1 Galvo loop was not working anymore. He tried to manually engaged it while the CITF was locked, but the loop opened immediately after.
Actions undertaken: We contacted the DET Oncall (Hui) that, with the support of Gouaty, found out that the related shutter was closed. After opening the shutter and re-centering the beam on the Quadrant they were able to close the loop again.

Images attached to this report
Comments to this report:
carbognani - 6:53 Friday 24 March 2023 (59402) Print this report

This night TCS HWS INJ measurements have completed correctly with 225 iteration under: /data/tcs/buffer/test/HWS-RC/INJ/20230324T0004

 

Search Help
×

Warning

Error

The present report has been modified outside this window. Please check for its integrity in the main page.

Refreshing this page will move this report into drafts.

×